OpenStudy (anonymous):

Consider the point (x,y) lying on the graph of the line 2x+4y=5. Let L be the distance from point (x,y) to the origin. Write L as a function of x

OpenStudy (anonymous):

This would involve using the distance formula. I can't remember the formula exactly. But that should help you solve it.

OpenStudy (anonymous):

d = √(x² + y²) Now, since (x,y) lies on the line 2x + 4y = 5 you can solve for y in terms of x to get y = (5-2x)/4 Now square y, y² = (5 - 2x)²/16 Substitue this into the formula for d above d = √(x² + (5 - 2x)²/16 ) Within the radical, combine over a common denominator of 16 d = √(16x² + (5 - 2x)² )/16 Factor out the radical 1/16, which is ¼, d = ¼ √(16x² + (5 - 2x)² ) Now, expand the binomial inside the radical and simplify, d = ¼ √(16x² + 25 - 20x + 4x² ) d = ¼ √(20x² - 20x + 25)